OpenStudy (annie12m):

Greg pays $0.10 per minute plus $5.00 per month for his phone bill. Write an Algebraic expression to show his total cost for X minutes of calls in one month. How much would Greg pay for one month if he uses his phone for 250 minutes? Show all work.

rebeccaxhawaii (rebeccaxhawaii):

0.10x+5.00=? x=250

rebeccaxhawaii (rebeccaxhawaii):

do u understand ?

OpenStudy (annie12m):

ya i do thanks

OpenStudy (annie12m):

like this??

OpenStudy (annie12m):

0.10x + 5 x = 250 0.10 (250) + 5 25 + 5 = 30

OpenStudy (annie12m):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

yes good job

OpenStudy (annie12m):

so greg pays $30 for that month?

rebeccaxhawaii (rebeccaxhawaii):

exactly

OpenStudy (annie12m):

ok thanks for your help!